# MLlm-DR: Towards Explainable Depression Recognition with MultiModal Large Language Models

### v1

@@ 6. Conclusion
In this paper, we propose a novel multimodal large language model (MLlm-DR). The model consists of a smaller LLMs and a lightweight query module (LQ-former), which are designed to generate explainable evaluation rationales and integrate multimodal data, respectively, enabling explainable and comprehensive depression diagnosis. This approach is closely aligned with clinical needs and holds significant practical application value. Our approach achieves state-of-the-art results on two interview-based benchmark datasets (CMDC and E-DAIC-WOZ), demonstrating its effectiveness and superiority. The construction of the training dataset is solely based on text, which may result in the loss of fine-grained emotional information from speech and visual cues, leading to potential bias. In future work, we hope creating a larger-scale, fine-grained depression label set for speech and visual data to further improve our research.

### v2

@@ 5.13. Analysis of Data Collection Methods
6. Limitations
Our model currently generates free-form textual descriptions to explain the patient’s mental state, which may raise potential ethical and privacy concerns (Zou et al., 2024; Liu et al., 2025; Wang and Ge, 2024; Hu et al., 2025). In future work, we plan to incorporate domain knowledge and contextual constraints to guide the explanation generation process, thereby improving the consistency and safety of explanations. For example, adopting template-based or semi-structured generation mechanisms can enhance the controllability and auditability of model explanations while maintaining readability. In addition, in multimodal fusion tasks, low-quality or distorted modalities may reduce the model’s ability to extract reliable features, leading to biased or unstable reasoning results (Sun et al., 2022b; Yu et al., 2025a). In future studies, we will explore integrating perceptual quality assessment (Yu et al., 2025b) and cross-modal enhancement mechanisms (Cheng et al., 2025) to improve the robustness and reasoning stability of the model.
@@ 7. Conclusion
In this paper, we propose a novel multimodal large language model (MLlm-DR). The model consists of a small-scale LLMs and a lightweight query module (LQ-former), which are designed to generate explainable evaluation rationales and integrate multimodal data, respectively, enabling explainable and comprehensive depression diagnosis. This approach is closely aligned with clinical needs and holds significant practical application value. Our approach achieves state-of-the-art results on two interview-based benchmark datasets (CMDC and E-DAIC-WOZ), demonstrating its effectiveness and superiority. Beyond performance gains, this work advances the integration of LLMs and multimodal affective computing. The unification of language-driven reasoning and multimodal perception in MLlm-DR provides an interpretable and extensible basis for trustworthy clinical decision support.
8. Acknowledgments
This work is supported by National Science and Technology Innovation 2030 Major Project under Grant No. 2022ZD0209103.